Complete the ordered pairs for the equation y = 5x + 5.
(?,0), (4/5,?), (0,?), (-4/5,?)

Answer by checkley75(3666)

MAYBE WE NEED TO START OFF WITH AN EXPLANATION OF WHAT YOU NEED TO DO.
ALL THAT IS REQURED TO SOLVE FOR THE OTHER UNKNOWN (X OR Y) IS TO SUBSTITUTE THE KNOWN VALUE AND SOLVE FOR THE OTHER UNKNOWN.
THE FIRST CHALLENGE IS (?,0) HERE WE HAVE Y=0 & NEED TO SOLVE FOR X.
SO WE REPLACE Y WITH 0 & SOLVE FOR X
Y=5X+5 OR 0=5X+5 OR 5X=-5 OR X=-5/5 OR X=-1 IS THE ANSWER FOR THE X VALUE(-1,0)
NOW WE HAVE A VALUE FOR THE X TERM 4/5. NOW REPLACE X BY 4/5 & SOLVE FOR Y.
Y=(4/5)5+5 OR Y=4+5 OR Y=9 IS THE ANSWER FOR THE Y TERM (4/5,9)
Y=5*0+5 OR Y=5 ANSWER (0,5)
Y=5(-4/5)+5 OR Y=-4+5 OR Y=1 ANSWER (-4/5,1)
---------------------------------------------------------
HOPE THIS ADDITIONAL INFORMATION HELPS.
IF NOT --- LET ME KNOW.